Let's say we want to withdraw $900 per month for 4 months. How much do we put into a savings account today that earns 6% compounded monthly? Assume we will have no money left in the account at the end of the 4 months.
Answer #1

Monthly nominal interest rate = 6%/12 = 0.5%

Savings account ($) = 900 x P/A(0.5%, 4) = 900 x 3.9505** = 3,555.45

**P/A(0.5%, 4) = [1 - (1.005)-4] / 0.005 = (1 - 0.9802) / 0.005 = 0.0198 / 0.005 = 3.9505